OpenStudy (anonymous):

Please help me 😔 10x^3 y^-5 z^-2 If x=3 y=2 z=5 Express your answer as a reduced fraction.

OpenStudy (mathstudent55):

First, replace each variable with its value. What do you get?

OpenStudy (anonymous):

10 3^3 2^-5 5^-2

OpenStudy (mathstudent55):

You're correct. \(10 \times (3^3) \times (2^{-5}) \times (5^{-2}) \) Now, turn each negative exponent into a fraction.

OpenStudy (mathstudent55):

We raise 3^3 and we change the negative exponents into fractions. \(10 \times (27) \times \dfrac{1}{2^5} \times \dfrac{1}{5^2} \) Now raise 2^5 and 5^2, and set the entire product as one single fraction.

OpenStudy (anonymous):

So i multiply 2^5 and 5^2

OpenStudy (mathstudent55):

First, raise the numbers to the exponents. Remember, 2^5 = 2 * 2 * 2 * 2 * 2, and 5^2 = 5 * 5

OpenStudy (anonymous):

So 10 and 25

OpenStudy (anonymous):

I dont get it

OpenStudy (mathstudent55):

No. Exponents mean repeated multiplication. 3^1 = 3 3^2 = 3 * 3 = 9 3^3 = 3 * 3 * 3 = 27 3^4 = 3 * 3 * 3 * 3 = 81 So 2^5 = 2 * 2 * 2 * 2 * 2 = 32 and 5^2 = 5 * 5 = 25

OpenStudy (mathstudent55):

\(=10 \times (27) \times \dfrac{1}{2^5} \times \dfrac{1}{5^2}\) \(=\dfrac{10 \times 27} {2^5 \times 5^2} \) \(=\dfrac{10 \times 27} {32 \times 25} \) \(=\dfrac{2 \times5 \times 27} {2 \times 16 \times 5 \times 5} \) \(=\dfrac{\cancel{2} \times \cancel{5} \times 27} {\cancel{2} \times 16 \times \cancel{5} \times 5} \) \(= \dfrac{27}{80} \)